OpenStudy (anonymous):

sin θ/2 = 0 find all solutions in the interval [0, 2pi]

OpenStudy (john_es):

\[\sin(\theta/2)=0\Rightarrow\theta/2=n\pi\Rightarrow \theta=n2\pi\ \ \ \ n=0,1,\ldots\]

OpenStudy (john_es):

There are only two solutions in the interval you want, n=0 and n=1.

OpenStudy (anonymous):

so are the solutions [ 0, 2pi] cause thats what i got at least

OpenStudy (john_es):

Yes, 0 and 2pi. Exact.

OpenStudy (john_es):

(not the interval, I mean)

OpenStudy (anonymous):

yah like when i did n=0 i got 2pi and when i did n=-1 i got 0 which were the only ones that fit the interval

OpenStudy (john_es):

Well, what I was intended to say is that the two only solutions in the interval are, \[\theta=0\]\[\theta=2\pi\]But may be we are saying the same.